OpenStudy (anonymous):

write a quadratic equation with the given roots. write the equation in the form ax^2+bx + c = 0 where a, b, and c are integers. 3/2, 1/4 A. 8x^2 - 14x + 3 = 0 B. 8x^2 - 3x + 14 = 0 C. 4x^2 - x + 6 = 0 D. 4x^2 - 6x + 1 = 0

OpenStudy (asnaseer):

if a quadratic equation has roots \(r_1\) and \(r_2\), then it can be written as:\[a(x-r_1)(x-r_2)=0\]if you expand this you will get:\[a(x^2-(r_1+r_2)x+r_1r_2)=0\]this should help you work out the answer

OpenStudy (anonymous):

i dont know what that is..

OpenStudy (asnaseer):

'a' in the equation above is some arbitrary constant - so you can ignore that for this problem - assume its equal to 1. so you are left with:\[x^2-(r_1+r_2)x+r_1r_2=0\]

OpenStudy (asnaseer):

just substitute the values given to you for \(r_1\) and \(r_2\).

OpenStudy (asnaseer):

in your case:\[r_1=\frac{3}{2}\]\[r_2=\frac{1}{4}\]

OpenStudy (asnaseer):

so what is the value of \(r_1+r_2\)?

OpenStudy (anonymous):

7/4

OpenStudy (asnaseer):

good, now what is the value of:\[r_1\times r_2\]

OpenStudy (anonymous):

3/8

OpenStudy (asnaseer):

great!, so just substitute these into this equation:\[x^2-(r_1+r_2)x+r_1r_2=0\]what do you end up with?

OpenStudy (anonymous):

x^2 - (7/4)x + 3/8 = 0

OpenStudy (asnaseer):

perfect! now we need to get rid of those nasty fraction. so just multiply everything by 8. what do you end up with?

OpenStudy (anonymous):

14.

OpenStudy (anonymous):

3

OpenStudy (asnaseer):

yes - so what does the final equation look like?

OpenStudy (asnaseer):

remember you needed to multiply EVERYTHING by 8, so: 8(x^2 - (7/4)x + 3/8) = 8*0

OpenStudy (asnaseer):

which expands to: 8x^2 - 8(7/4)x + 8(3/8) = 0

OpenStudy (anonymous):

8x^2 - (14)x+(3) = 0

OpenStudy (asnaseer):

thats it! great work! :)

OpenStudy (anonymous):

whats r?

OpenStudy (asnaseer):

where do you see r?

OpenStudy (asnaseer):

do you mean \(r_1\) and \(r_2\)?

OpenStudy (anonymous):

its the answer. r = 36, r = 12 r = 51 r = 13

OpenStudy (asnaseer):

what is the full question?

OpenStudy (anonymous):

oh nvm.

OpenStudy (asnaseer):

ok :)

OpenStudy (anonymous):

thansk

OpenStudy (asnaseer):

yw :)